OpenStudy (anonymous):

A bowl contains 4 yellow marbles and 3 red marbles. Eric randomly draws 3 marbles from the bowl. He replaces each marble after it is drawn. What is the probability that he draws 1 yellow marble and 2 red marbles?

OpenStudy (anonymous):

solve the equation 4/7 x 3/7 x 3/7 (x=multiply)

OpenStudy (michele_laino):

please you have to apply the definition of probability, namely probability is equal to the ratio between favorable cases over possible cases, of course both favorable cases and possible cases are equally probable.

OpenStudy (michele_laino):

for example probability of getting 1 red marble is: 3/7 and probability of getting yellow marble is 4/7

OpenStudy (michele_laino):

now, keep in mind that probability of getting 1 yellow marble and 1 red marble and 1 red Marble is the product of the single probability, since after 1 draw the marble is replaced in the bowl

OpenStudy (anonymous):

exactly so that was the correct equation

OpenStudy (michele_laino):

yes, I think so!

OpenStudy (anonymous):

so out of these answers, would it be A? A) 36/343 B) 108/343 C) 12/35 D) 4/35 @Michele_Laino @Jrigby20

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

@kianatalani

OpenStudy (michele_laino):

we have to compute this: \[\frac{ 4 }{ 7 }*\frac{ 3 }{ 7 }*\frac{ 3 }{ 7 }=...\] please continue

OpenStudy (anonymous):

i already did, a is correct

OpenStudy (anonymous):

4*3*3=36

OpenStudy (anonymous):

and 7*7*7=343

OpenStudy (anonymous):

thus 36/343

OpenStudy (anonymous):

@kianatalani @Michele_Laino hey kianatalani, which way was easier for you to understand?

OpenStudy (anonymous):

your way :) thanks for the help! @Jrigby20

OpenStudy (anonymous):

np kiantalani, glad i could help